# GPT4Scene: Understand 3D Scenes from Videos with Vision-Language Models

### v2

@@ 6 Conclusion
We introduce GPT4Scene, a framework that enhances Vision-Language Models (VLMs) to comprehend 3D scenes directly from pure vision input. Our pilot study reveals that global scene information and the correspondence between video frames and objects in the global context are essential for improving the 3D understanding of VLMs. We propose reconstructing a 3D point cloud from input video to generate a BEV (Bird’s Eye View) image for global information. We establish global and local data correspondences by adding spatial-temporal object markers (STO markers) to both video frames and BEV images. GPT4Scene achieves excellent performance with closed-source VLMs like GPT-4o using zero-shot inference. For smaller VLMs, such as Qwen2-VL, we create the ScanAlign dataset, which includes video frames, BEV images, STO markers, and 165K text annotations. After fine-tuning, VLMs show significant improvements in scene understanding, reaching state-of-the-art performance in question-answering tasks. Moreover, fine-tuned VLMs can perform well on QA tasks with only original video frames, demonstrating that GPT4Scene enables VLMs to understand 3D scenes effectively.
7 Appendix
@@ Appendix C Qualitative Results
Figures 6, 7, 8 and 9 presents our qualitative results, derived from Qwen2-VL after fine-tuning on ScanAlign. Figures 6 and 7 showcase the results of 3D question answering, where only unmarked videos were provided as input. We observe that the model can answer questions from ScanQA [6] and SQA3D [66]. Meanwhile, Figures 8 and 9 illustrate the model’s performance on 3D dense captioning and 3D visual grounding tasks, where marked inputs were required. These results demonstrate that our model can provide accurate responses, relying solely on visual information without requiring 3D point cloud input.
@@ Appendix D Full Quantitive Results
Here, we present the complete metrics for all five benchmarks. Table 9 and Table 10 show results for ScanQA [6] and SQA3D [66]. Table 11 provides the full results for Scan2Cap [22], while Table 12 and Table 13 present the results for ScanRefer [13] and Multi3DRef [114]. Our model significantly improves across all benchmarks, highlighting that only pure vision language models can understand 3D scenes effectively.

### v4

@@ 5 Conclusion
We propose GPT4Scene, a framework enhancing Vision-Language Models (VLMs) to understand 3D scenes directly from visual inputs. By reconstructing 3D point clouds for Bird’s Eye View (BEV) images and aligning video frames with spatial-temporal object (STO) markers, we enable global-local scene comprehension. GPT4Scene achieves state-of-the-art 3D QA performance with zero-shot GPT-4o and fine-tuned smaller VLMs (e.g., Qwen2-VL) using our ScanAlign dataset. Fine-tuned models even excel with raw video inputs, proving effective 3D understanding.
@@ Appendix C Qualitative Results
Figures 5, 6, 7 and 8 present our qualitative results obtained from Qwen2-VL after fine-tuning on ScanAlign. Figures 5 and 6 demonstrate 3D question answering performance using only unannotated video inputs as context. Our observations indicate the model’s capability to effectively address queries from both ScanQA [5] and SQA3D [66] benchmarks. Meanwhile, Figures 7 and 8 exhibit the model’s competence in 3D dense captioning and 3D visual grounding tasks requiring annotated inputs. These outcomes substantiate that our framework can generate precise responses through visual information processing alone, without dependency on 3D point cloud inputs. Here we emphasize that post-training with the GPT4Scene framework, the model achieves accurate QA performance using only pure video input, thereby demonstrating GPT4Scene’s effectiveness in enhancing visual comprehension capabilities.
@@ Appendix D Full Quantitive Results
Here, we present the complete metrics for all five benchmarks. Table 6(c) and Table 13 show results for ScanQA [5] and SQA3D [66]. Table 14 provides the full results for Scan2Cap [21], while Table 15 and Table 16 present the results for ScanRefer [12] and Multi3DRef [121]. Our model significantly improves across all benchmarks, highlighting that only pure vision language models can understand 3D scenes effectively.